OpenStudy (toman):

The ratio of the ages of a father and son is 17:7 respectively. Six years ago the ratio of their ages was 3:1 respectively. What is the son's present age.

OpenStudy (anonymous):

dad's age now / son's age now = 17/7 dad's age 6 yrs ago / son's age 6 yrs ago = 3/1

OpenStudy (welshfella):

let x = fathers age and y = sons age now x/y = 17/7 giving x = 17y/7................(1) also (x -6) / (y - 6) = 3/1 giving x - 6 = 3(y-6) .......................(2) solve equations (1) and (2) to find x and y

OpenStudy (toman):

Thank you. c:

OpenStudy (anonymous):

Sounds good